OpenStudy (anonymous):

Polynomials find all non-Zero polynomials P(x) with real coefficients such that: \[P(2x)=P'(x) P''(x)\]

OpenStudy (zarkon):

I only see one non-constant polly that works

OpenStudy (anonymous):

what is that?

OpenStudy (zarkon):

\[p(x)=\frac{4}{9}x^3\]

OpenStudy (turingtest):

how?

OpenStudy (anonymous):

let's suppose that the degree of P(x) is n hence degree of P(2x) is : n degree of P'(x) is : n-1 degree of P''(x) is : n-2 degree of P'(x) P''(x) is : n-1+n-2=2n-3 according to the equation degree of P(2x) is equal to degree of P'(x) P''(x) so n=2n-3 --> n=3 then we have \[P(x)=ax^3+bx^2+cx+d \\ P(2x)=8ax^3+4bx^2+2cx+d \\ P'(x)=3ax^2+2bx+c\\ P''(x)=6ax+2b\] according to the equation \[ 8ax^3+4bx^2+2cx+d =(3ax^2+2bx+c)(6ax+2b)\] simplifying and solve for coefficients gives a=4/9 and b=c=d=0 then the only answer is \[P(x)=\frac{4}{9} x^3\]

OpenStudy (turingtest):

nice!!!

OpenStudy (anonymous):

why do you choose \[P(x)=ax^3+bx^2+cx+d \]

OpenStudy (anonymous):

because we found the degree of polynomial and that was n=3

OpenStudy (anonymous):

oh I see now..